### v1

# Proton mass decompositions in the NNLO QCD

Abstract
Proton matrix elements of the QCD energy-momentum tensor are expressed by the gravitational form factors. The forward values of the gravitational form factors allow for a decomposition of the proton mass into the contributions from quarks and gluons, and the further subdivisions into the contributions from quark masses and from the QCD trace anomaly may be considered. We present the most recent evaluations of these mass decompositions, using a recent quantitative evaluation of the forward values of relevant gravitational form factors at the next-to-next-to-leading order (NNLO) QCD. We also calculate the renormalization scale dependence of each component within these decompositions. The results indicate several shortcomings of those decompositions. We propose a new decomposition according to the twist of the relevant operators, i.e., classifying the effects via the partonic motion according to asymptotic freedom and the parton correlation by non-perturbative QCD interactions. We demonstrate the improvements by our new decomposition. We also present the results for the pion, which exhibit quite different parton-correlation behaviors from the proton.
The energy-momentum tensor (EMT) plays fundamental roles in quantum field theories, arising in the generators for the Poincaré, dilatation, and conformal transformations [1, 2]. In QCD the EMT is composed of several composite operators, and their physical interpretations are not straightforward due to their sophisticated renormalization properties [3, 4, 5] and the associated trace anomalies [6, 7, 8, 9, 10]. These complications, in particular, have been preventing the attempts [11, 12, 13, 14, 15, 16, 17] to understand the mass structure of the proton from the EMT.
The symmetric EMT in QCD, $T^{\alpha\beta}=T^{\alpha\beta}_{q}+T^{\alpha\beta}_{g}$, reads [11, 2, 18]
$T^{\alpha\beta}_{q}=i\bar{\psi}\gamma^{(\alpha}\overleftrightarrow{D}^{\beta)}\psi,\;\;\;T_{g}^{\alpha\beta}=-F^{\alpha\lambda}F^{\beta}_{\ \lambda}+\frac{\eta^{\alpha\beta}}{4}F^{2},$ (1)
where $T^{\alpha\beta}_{q}$ and $T_{g}^{\alpha\beta}$ are the gauge-invariant quark part and gluon part, $D^{\alpha}=\partial^{\alpha}+igA^{\alpha}$, $\overleftrightarrow{D}^{\alpha}\equiv(\overrightarrow{D}^{\alpha}-\overleftarrow{D}^{\alpha})/2$, $R^{(\alpha}S^{\beta)}\equiv(R^{\alpha}S^{\beta}+R^{\beta}S^{\alpha})/2$, $F^{\alpha\lambda}$ is the gluon field strength tensor, and $\eta^{\alpha\beta}$ is the metric tensor; we have neglected the gauge-variant terms, i.e. the ghost terms and the gauge fixing term, as they do not affect our final results  [^1] The hadron matrix elements of those gauge-variant terms vanish, see e.g., [3, 19].. $T^{\alpha\beta}$ can be constructed from the canonical EMT by adding appropriate total-divergence contributions to make it symmetric, $T^{\alpha\beta}=T^{\beta\alpha}$, while preserving the continuity equation, $\partial_{\alpha}T^{\alpha\beta}=0$. All composite operators presented in this Letter are renormalized in the MS-like (MS, $\overline{\rm MS}$) schemes in the dimensional regularization [3], and we follow the conventions of [20].
Our proton state $|p\rangle$ has the eigenvalue $p^{\alpha}$ for the 4-momentum operator defined as the volume integral of $T^{0\alpha}$ and is normalized as $\langle p^{\prime}|p\rangle=2p^{0}(2\pi)^{3}\delta^{(3)}(p^{\prime}-p)$. Then, translation invariance allows us to show $\langle p|T^{\alpha\beta}|p\rangle=2p^{\alpha}p^{\beta}$, from which the well-known expressions for the proton mass $M$ read,
$2M^{2}=\left.\langle p|T^{00}|p\rangle\right|_{\bm{p}=\bm{0}},\;\;\;\;\;\;2M^{2}=\langle p|\eta_{\alpha\beta}T^{\alpha\beta}|p\rangle,$ (2)
as the “rest energy” and as the “invariant mass”.
The invariant mass of (2) leads to the well-known proton mass formula [21] associated with the trace anomaly [6, 7, 8], i.e., $T^{\alpha}_{\ \alpha}=\frac{\beta(g)}{2g}F^{2}+[1+\gamma_{m}(g)]m\bar{\psi}\psi$, with the QCD $\beta$-function $\beta(g)$ and the anomalous dimension for the quark mass, $\gamma_{m}(g)$. Thus, the proton mass is expressed by the non-perturbative matrix elements $\langle p|F^{2}|p\rangle$, $\langle p|m\bar{\psi}\psi|p\rangle$; these matrix elements of the Lorentz scalar operators equal the corresponding matrix elements at the rest frame, as $\langle p|{\cal O}(0)|p\rangle$ $\langle p|{\cal O}(0)|p\rangle|_{\bm{p}=\bm{0}}=2M\langle\int d^{3}x{\cal O}(x)\rangle$, with ${\cal O}(x)=F^{2}(x),m\bar{\psi}\psi(x)$, where we have introduced the shorthand notation for the expectation value in the rest frame as $\langle\,\cdots\,\rangle\equiv(\langle p|\cdots|p\rangle/\langle p|p\rangle)_{\bm{p}=\bm{0}}$. As discussed below, $\langle p|m\bar{\psi}\psi|p\rangle$ is expressed by the sigma terms determined (semi)emprically; then, the condition that the second equation of (2) should reproduce the value of the proton mass allows us to evaluate the proton matrix element $\langle p|F^{2}|p\rangle$. Now, using those non-perturbative values of $\langle p|F^{2}|p\rangle$ as well as of $\langle p|m\bar{\psi}\psi|p\rangle$, the separation of $T^{\alpha\beta}$ into (1) allows us to decompose the proton mass into the quark and gluon contributons as $M={\cal M}_{q}+{\cal M}_{g}$ (see (14) below), combined with the trace anomaly, as taking trace after renormalization  [^2] Taking trace and performing renormalization do not commute; if their order were reversed, as taking trace before renormalization, the RHS of (3) would be modified into $m\bar{\psi}\psi$, which corresponds to the “D2 scheme”[14, 17, 15]., for each part of (1),
$T_{q\ \alpha}^{\alpha}=m\bar{\psi}\psi+\frac{{{\alpha_{s}}}}{{4\pi}}\left(\frac{{{n_{f}}}}{3}F^{2}+\frac{{4{C_{F}}}}{3}m\bar{\psi}\psi\right)+\cdots,$ (3)
and $T_{g\ \alpha}^{\alpha}=T^{\alpha}_{\ \alpha}-T_{q\ \alpha}^{\alpha}$, with $\alpha_{s}\equiv g^{2}/4\pi$, for $N_{c}$ color with $C_{F}=(N_{c}^{2}-1)/2N_{c}$; (3) holds in the MS-like schemes, and the ellipses stand for the corrections at the two loops (${\cal O}(\alpha_{s}^{2})$) and higher, which are derived in [5, 9, 10]. The quantitative results using the three-loop formulas [9] for (3) will be discussed in Figs. 2, 4 below.
The rest energy of (2) reads $M=(\langle p|T^{00}|p\rangle/2M)_{\bm{p}=\bm{0}}$$\langle\,H\,\rangle$, and $H=\int d^{3}x[T_{q}^{00}(x)+T_{g}^{00}(x)]$, i.e.,
$H=\int d^{3}x\Bigl[\bar{\psi}\bigl(i\bm{\gamma}\cdot\overleftrightarrow{\bm{D}}+m\bigr)\psi+\frac{1}{2}\bigl(\bm{E}^{2}+\bm{B}^{2}\bigr)\Bigr],$ (4)
denotes the QCD Hamiltonian composed of the quark Dirac Hamiltonian and the chromoelectric/chromomagnetic Hamiltonian, as derived from (1), so that $M=\langle\,H\,\rangle$ leads to the mass formula as the sum of the rest energies, $M=U_{q}+U_{g}$, due to the quark and gluonic degrees of freedom. The corresponding expectation values of the quark and gluon Hamiltonian operators of (4) could be computed by non-perturbative methods like lattice QCD [22]. Here, alternatively, we represent and evaluate them using the gravitational form factors, $A_{r}(\Delta^{2},\mu)$ and $\bar{C}_{r}(\Delta^{2},\mu)$ with $r=q,g$, which are the functions of $\Delta^{2}\equiv(p^{\prime}-p)^{2}$ as well as the renormalization scale $\mu$, and are defined for each of (1) as
$\!\!\!\langle p^{\prime}|T_{r}^{\alpha\beta}|p\rangle=\bar{u}(p^{\prime})\bigl[A_{r}\gamma^{(\alpha}\overline{p}^{\beta)}+\bar{C}_{r}M\eta^{\alpha\beta}\bigr]u(p)+\cdots,$ (5)
where $\overline{p}^{\alpha}\equiv(p^{\alpha}+p^{\prime\alpha})/2$, $u(p)$ is the proton spinor, and the ellipses stand for the terms associated with the other Lorentz structures that vanish as $\Delta=p^{\prime}-p\to 0$ (see e.g., [18, 23, 24, 25]). The sum rule, $\bar{C}_{q}(\Delta^{2},\mu)+\bar{C}_{g}(\Delta^{2},\mu)=0$, is derived by contracting (5) with $\Delta_{\alpha}$ and using $\partial_{\alpha}T^{\alpha\beta}=0$. Also, $\langle p|T^{\alpha\beta}|p\rangle=2p^{\alpha}p^{\beta}$ ensures the forward sum rule, $A_{q}(0,\mu)+A_{g}(0,\mu)=1$. Physically, the gravitational form factors represent the coupling of the proton with a graviton. Instead of probing them via such impractical graviton-proton processes, the gravitational form factors can be accessed through the spin-2 contributions induced by the two vector currents; e.g., with the two electromagnetic currents, the deeply virtual Compton scatterings [26] in JLab [27] and the future electron-ion collider (EIC) allow us to probe the gravitational form factors. Substituting the forward limit of (5) into $M=\langle\,H\,\rangle$, we obtain the “two-term decomposition” [13] using the $\Delta=0$ values of the gravitational form factors as
$M=U_{q}+U_{g},\;\;\;\;\;U_{r}=M\left(A_{r}(0,\mu)+\bar{C}_{r}(0,\mu)\right).$ (6)
Recently, the values of $\bar{C}_{r}(0,\mu)$, as well as of $A_{r}(0,\mu)$, as a function of $\mu$ have become available in the next-to-next-to-leading order (NNLO), at the level of accuracy with the $\rm uncertainty$$\lesssim$ a few percent. It is known [18, 23] that the matrix elements (5) for $A_{r}(0,\mu)$ coincide with the first moments of the proton’s parton distribution functions (PDFs) at the scale $\mu$ as $A_{q}(0,\mu)=\sum_{f}\int_{0}^{1}dxx(q_{f}(x,\mu)+q_{\bar{f}}(x,\mu))$, summed over the quark and antiquark distributions of flavor $f=u,d,\ldots$, and similarly for $A_{g}(0,\mu)$ using the gluon distribution. The $\mu$-dependence of $A_{r}(0,\mu)$ is governed by the renormalization group (RG) equations for the operators (1), and these coincide with the first moment of the evolution equations for the PDFs. The explicit NNLO formulas of $A_{q}(0,\mu)$ in the $\overline{\rm MS}$ scheme as the solution of the corresponding three-loop RG equations are given in e.g., [20], containing one single nonperturbative parameter as the value at a starting scale of the RG evolution; e.g., $A_{q}\left(0,\mu_{i}\right)=0.613$ at the starting scale $\mu_{i}\equiv 1.3~{\rm GeV}$ is determined in the $\overline{\rm MS}$ scheme by the NNLO global QCD analysis of the experimental data for the nucleon’s PDFs (CT18 [28]), representing the total momentum fraction carried by the three flavors, $u,d,s$, which are the active quark flavors at the starting scale $\mu_{i}$. The uncertainties for $A_{r}(0,\mu)$ as the NNLO solution are of a few percent level. For $\mu_{0}\equiv 2$ GeV, the CT18 global analysis with the variable flavor number [28] gives $\sum_{f=u,d,s}\int_{0}^{1}dxx[q^{\rm CT18}_{f}(x,\mu_{0})+q^{\rm CT18}_{\bar{f}}(x,\mu_{0})]=0.576(11)$, $A_{g}^{\rm CT18}(0,\mu_{0})=0.414(8)$, while the above input value, $A_{q}(0,\mu_{i})=0.613$, is evolved to $\left.A_{q}(0,\mu_{0})\right|_{{n_{f}}=3}=0.574$, $\left.A_{g}(0,\mu_{0})\right|_{{n_{f}}=3}=0.426$, in the NNLO with the fixed flavor number $n_{f}=3$; thus, in the following evaluations, we shall use,
$\!\!\!\!\left.A_{q}\left(0,\mu_{0}\right)\right|_{{n_{f}}=3}=0.57(1),\ \left.A_{g}\left(0,\mu_{0}\right)\right|_{{n_{f}}=3}=0.43(1).$ (7)
The values of $\bar{C}_{r}(0,\mu)$ are also available with the same level accuracy as in $A_{r}(0,\mu)$: the trace of both sides of (5) allows us to write down the formula [5],
$\bar{C}_{r}(0,\mu)=-\frac{1}{4}A_{r}(0,\mu)+\frac{1}{8M^{2}}\langle p|T_{r\ \alpha}^{\alpha}|p\rangle,$ (8)
and both terms in the RHS are evaluated at the NNLO accuracy [20], substituting the above-mentioned three-loop evaluation of $A_{r}(0,\mu)$ and of (3). The explicit NNLO formula [20] of $\bar{C}_{r}(0,\mu)$ in the $\overline{\rm MS}$ scheme gives [^3] We note that ${\bar{C}_{q}(0,\mu_{0})}|_{n_{f}=3}=-0.163(3)$ presented in [20] shows the uncertainty due to those of the sigma terms $\sigma_{\pi N},\sigma_{s}$, while (8) tells us that it should be combined with the uncertainty of (7). This results in (9).
$\left.{\bar{C}_{q}(0,\mu_{0})}\right|_{{n_{f}}=3}=-\left.{\bar{C}_{g}(0,\mu_{0})}\right|_{{n_{f}}=3}=-0.163(4).$ (9)
Those NNLO values (7),(9) give $M=[0.41(1)+0.59(1)]M$ at the scale $\mu=\mu_{0}$ (=2=2 GeV), using (6).
This result may be elaborated by separating the contribution due to the quark mass term in the Dirac Hamiltonian of (4), so that the “three-term decomposition” is obtained as $M=(U_{q}-M_{m})+U_{g}+M_{m}$ [14, 17],with
$M_{m}\equiv\bigl\langle\,\int d^{3}xm\bar{\psi}\psi\,\bigr\rangle=\frac{\left.\langle p|m\bar{\psi}\psi|p\rangle\right|_{\bm{p}=\bm{0}}}{2M}.$ (10)
This may be elaborated further, separating each of (1) into the traceless and trace parts as, $T_{r}^{\alpha\beta}=(T_{r}^{\alpha\beta}-\eta^{\alpha\beta}T_{r\,\lambda}^{\lambda}/4)+\eta^{\alpha\beta}T_{r\,\lambda}^{\lambda}/4$, with the latter being related to the trace-anomaly contribution,
$M_{a}\equiv({\cal M}_{q}-M_{m}+{\cal M}_{g})/4=(M-M_{m})/4$
$=\frac{1}{4}\bigl\langle\ \int d^{3}x\bigl[\ \frac{\beta(g)}{2g}F^{2}+\gamma_{m}(g)m\bar{\psi}\psi\ \bigr]\ \bigr\rangle,$ (11)
so that
$M=M_{q}+M_{g}+M_{m}+M_{a},$ (12)
with $M_{q}=U_{q}-M_{m}-({\cal M}_{q}-M_{m})/4$, $M_{g}=U_{g}-{\cal M}_{g}/4$; this corresponds to the famous “four-term decomposition” proposed by Ji [11, 4, 12] and coincides with Eq.(62) of [17]. For the $n_{f}=3$ active quark flavors, $M_{m}$ of (10) can be expressed as $M_{m}=\langle p|[(m_{u}+m_{d})/2](\bar{u}u+\bar{d}d)+m_{s}\bar{s}s|p\rangle/2M=\sigma_{\pi N}+\sigma_{s}$, up to small isospin-violating corrections of ${\cal O}(m_{d}-m_{u})$, using the pion-nucleon sigma-term $\sigma_{\pi N}$ and the strangeness content $\sigma_{s}$. Here, we use $\sigma_{\pi N}=0.0591(35)$ GeV due to a recent phenomenological analysis [29] and $\sigma_{s}=0.0456(62)$ GeV due to a recent lattice QCD determination [30]; these non-perturbative matrix elements arise also in (8) with (3), and the above-mentioned NNLO values of $\bar{C}_{q,g}(0,\mu)$ (e.g., (9)) are calculated using the same input values for $\sigma_{\pi N},\sigma_{s}$.
Our result for (12) using the above-explained non-perturbative inputs provides the NNLO update for the evaluation of Ji’s four-term decomposition of the proton mass, as shown in Fig. 1(a).
Figure 1: The ratios $M_{j}/M$ ($j=q,g,m,a$) in Ji’s four-term decomposition of the proton mass, (12), in the ($\overline{\rm MS}$) NNLO QCD with the $n_{f}=3$ active flavors: (a) at $\mu=2$ GeV; (b) as functions of $\mu$ with the dashed line indicating the case (a).
Our numerical calculations are performed for the fixed number $n_{f}=3$ active quark flavors using the three-loop $\overline{\rm MS}$ running coupling for $\alpha_{s}(\mu)$  [^4] We determine $\alpha_{s}(\mu)$ as the exact numerical solution [31] of the corresponding RG equation controlled by the three-loop $\beta$ function. We use $\Lambda_{\rm QCD}^{(3)}\simeq 0.3359$ GeV, i.e., $\alpha_{s}(\mu=1~\rm{GeV})\simeq 0.4736$, as in [20].. Fig. 1(b) shows the $\mu$ dependences of each of four terms in (12), i.e., the results of the NNLO QCD evolution from each contribution of Fig. 1(a): $M_{q}$ and $M_{g}$ exhibit the significant $\mu$ dependence with their crossing around the hadronic scale, while $M_{m}$ as well as $M_{a}$ is RG invariant, see (10), (11). Those NNLO results for the proton are accurate to a few percent uncertainties; e.g., the numbers presented inside the pie chart of Fig. 1(a) have the uncertainties at most $\pm 1$ only in the last digit, and are consistent with the corresponding lattice QCD calculation [22] of $M=\langle\,H\,\rangle$. It is remarkable that the present NNLO QCD results with a few (semi)empirical non-perturbative inputs have the uncertainties smaller considerably than those in the lattice QCD results.
Our NNLO evaluation of (12) for the proton mass exhibits several shortcomings  [^5] Similar shortcomings (i)-(iii) arise also in the three-term decomposition discussed above (10); e.g., (ii) was pointed out in [14].:
(i) Strong $\mu$ dependence: one has to be careful for the strong $\mu$ dependence in Fig. 1(b) when referring to the pie chart of Fig. 1(a), as the latter would not represent a universal picture over the relevant scales.
(ii) Potential heavy-quark contribution: the decoupling of the quarks heavier than the $s$ quark, for the quark mass term (10), is not manifest; e.g., the inclusion of the charm content, $\sigma_{c}=\langle p|m_{c}\bar{c}c|p\rangle/2M=0.107(22)~{\rm GeV}$, which is obtained by recent lattice QCD calculation [30] (see also [32, 33]), would double the contribution of $M_{m}$, leading to its $\sim 100$% uncertainty, and also cause the corresponding modifications in $M_{q}$ as well as $M_{a}$.
(iii) Incomplete twist separation: using the gravitational form factors of (5), each term of (12) is expressed [11, 4] as $M_{q}=3MA_{q}(0,\mu)/4-3M_{m}/4$, $M_{g}=3MA_{g}(0,\mu)/4$, and $M_{a}=M/4-M_{m}/4$, where $M_{m}$ of (10) is a twist-four matrix element. Since $A_{r}(0,\mu)$ is of twist two, as convinced shortly, $M_{g}$ ($M_{a}$) is of twist two (four), while $M_{q}$ is a mixture of the twist-two and twist-four quantities; (12) is not fully organized according to the twist.
The composite operators are classified by the twist that equals their “dimension minus spin”, and the traceless and trace parts of (1) are of twist two and four, respectively. Comparing this with (5), $A_{r}(0,\mu)$ is of twist two, and this implies (iii) above. The studies of hard processes like the deep inelastic lepton-proton scatterings (DIS) clarified [19] that the twist-two operators in QCD are associated with the independent partonic motion inside the proton, according to the asymptotic freedom, while the non-perturbative QCD interactions induce the correlations among the partons, emerging as matrix elements of the operators of twist four and higher  [^6] For the quantities dependent on the polarization, the twist-three operators also play roles to describe the spin-dependent partonic correlations, see e.g., [19, 34, 35, 36].. As noted in (iii), $M_{q}$ does not match with such physical meaning of the twist, although $M_{q}$ is intended [11, 4] to represent the contribution of the quark’s gauge-invariant “kinetic and potential energies” operator $\bar{\psi}i\bm{\gamma}\cdot\overleftrightarrow{\bm{D}}\psi$ of (4). The physical meaning of the twist calls for the rearrangement of the conventional decomposition (12), such that the terms of different twist in each of the quark and gluon contributions are completely separated (see the formulas in (iii)): $M_{q}+M_{m}=(M_{q}+3M_{m}/4)+M_{m}/4=3MA_{q}(0,\mu)/4+M_{m}/4$, where $M_{m}/4$ is to be combined with (11), leads to the new four-term decomposition,
$\!M=\frac{3}{4}MA_{q}(0,\mu)+\frac{3}{4}MA_{g}(0,\mu)+\frac{1}{4}{\cal M}_{q}+\frac{1}{4}{\cal M}_{g},$ (13)
where the last two terms of twist four are expressed as
${\cal M}_{r}=\frac{\langle p|T_{r\ \alpha}^{\alpha}|p\rangle}{2M}=M\left(A_{r}(0,\mu)+4\bar{C}_{r}(0,\mu)\right),$ (14)
using (8).
Figure 2: Each term, divided by $M$, of the new four-term decomposition of the proton mass, (13), in the ($\overline{\rm MS}$) NNLO QCD with the $n_{f}=3$ active flavors: (a) at $\mu=2$ GeV; (b) as functions of $\mu$ with the dashed line indicating the case (a).
Remarkably, this resolves not only (iii) but also (i), (ii), as we now demonstrate: using the NNLO values (7), (9) and their evolution as functions of $\mu$, we evaluate each term of (13) as displayed in Fig. 2, which is accurate to a few percent uncertainties, similarly as in Fig. 1; the numbers shown inside Fig. 2(a) have the uncertainties at most $\pm 1$ only in the last digit. Note that the twist-four quark contribution ${\cal M}_{q}/4$ of (13) is too small (2-2% of $M$) to show up in Fig. 2(a). In Fig. 2(b), we see that the $\mu$ dependence of the twist-two terms of (13) yields the similar values $3A_{q}(0,\mu)/4\simeq 3A_{g}(0,\mu)/4\simeq 0.37$ for $10~{\rm GeV}\lesssim\mu\lesssim 20~{\rm GeV}$  [^7] The approach to the well-known asymptotic limit, $A_{q}(0,\infty)=n_{f}/(4C_{F}+n_{f})$, $A_{g}(0,\infty)=4C_{F}/(4C_{F}+n_{f})$, is quite slow, see, e.g., [20].; hence, the crossing of the quark and gluon twist-two contributions around the hadronic scale does not arise in Fig. 2(b), and (i) is resolved. Now, Fig. 2(a) could represent a universal qualitative picture of the proton-mass structure.
As explained above (13), the ${\cal O}(\alpha_{s}^{0})$ contribution associated with the quark mass operator in the rest energy of (2) is $M_{m}/4$, and it is organized in (13) as a twist-four term in ${\cal M}_{q}/4$, see (3). When $\sigma_{Q}=\langle p|m_{Q}\bar{Q}Q|p\rangle/2M$ of (ii) due to a heavy quark $Q$ is included by the replacements, $M_{m}/4\to(M_{m}+\sigma_{Q})/4$ and $n_{f}\to n_{f}+1$ in ${\cal M}_{q}/4$ of (14) with (3), the cancellation of the leading terms of those new contributions is ensured by the heavy-quark expansion of the corresponding mass operator,
$\langle h|m_{Q}\bar{Q}Q|h\rangle=-\frac{1}{12\pi}\langle h|\alpha_{s}F^{2}|h\rangle+\cdots\ ,$ (15)
which was derived in [21, 37, 38] for a state $|h\rangle$ with no valence heavy $Q$, calculating the triangle-type heavy-$Q$ loop diagrams that induce the long-distance gluon fields, and the ellipses denote the corrections suppressed by the two or more inverse-powers of $m_{Q}$. Therefore, the decoupling of a heavy flavor $Q$ is manifest for the combination of the operators in (3) relevant to ${\cal M}_{q}/4$ of (13), using (15); this logic may be extended to higher orders by the help of the formulation of [39]. Now (ii) is resolved.
The present framework can be applied also to the pion case, replacing the non-perturbative inputs by the corresponding pion quantities, which we denote with the superscript $\pi$”.
Figure 3: Same as Fig. 1, but for the pion case.
As explained in [20], the global QCD analyses for the pion’s PDFs are available at the NLO level, and the resultant value, $A_{q}^{\pi}(0,\mu_{0})=0.58(9)$ [40], has rather large ($\sim$ ten percent level) uncertainties (see also [41, 42, 43]), while chiral perturbation theory gives the accurate value for $M_{m}^{\pi}=\langle\pi|m\bar{\psi}\psi|\pi\rangle/2M^{\pi}$ corresponding to (10): $M_{m}^{\pi}=M^{\pi}/2$, to the $O(6\%)$ correction [44, 45]. Using these input values, $A_{q,g}^{\pi}(0,\mu)$, $\bar{C}^{\pi}_{q,g}(0,\mu)$ have been calculated [20] by our NNLO formulas in the $\overline{\rm MS}$ scheme as, e.g., $\bar{C}^{\pi}_{q}(0,\mu_{0})|_{{n_{f}}=3}=-\bar{C}^{\pi}_{g}(0,\mu_{0})|_{{n_{f}}=3}=-0.03(2)$ with $\mu_{0}=2$ GeV. These values allow us to evaluate the pion-mass decompositions corresponding to the four-term formulas (12) and (13), as displayed in Figs. 3 and 4, respectively, where the presented results have the uncertainties of $\sim$ ten percent level; we recognize that, also in the pion case, the points (i)-(iii) discussed above are improved using (13). The comparison between Figs. 2 and 4 reveals: the twist-two contributions behave similarly, while the proton and the pion are discriminated by the behaviors of the twist-four components, as driven by the large quark-mass term $M_{m}^{\pi}$ for the pion, which reflects the Nambu-Goldstone boson nature. Such difference could not be observed, if the last two terms of twist four in (13) were not subdivided into the quark and gluon contributions, because ${\cal M}_{q}/4+{\cal M}_{g}/4=M/4$; on the other hand, such difference could be observed, when the “D2 scheme” [14, 17] is employed as in [15] for the twist-four contributions in (13) by making the replacements ${\cal M}_{q}\to M_{m}$, ${\cal M}_{g}\to M_{a}$, but we stick to the $\overline{\rm MS}$ scheme to treat the twist-four effects as well as twist-two effects for our quantitative evaluation.
Figure 4: Same as Fig. 2, but for the pion case.
The twist-four contributions in Figs. 24 have the very weak $\mu$ dependence, and represent also the quark/gluon decomposition [9], $M={\cal M}_{q}+{\cal M}_{g}$, of the invariant mass of (2). This decomposition, composed of the twist-four effects only, coincides in total with the rest-mass decomposition (13), involving also the twist-two effects; this is guaranteed by the above-mentioned sum rule, $\bar{C}_{q}+\bar{C}_{g}=0$, compare (13) and (14). Apparently, the different characteristics arising in those different types of decompositions are caused by the behavior of $\bar{C}_{q}(0,\mu)$ ($\bar{C}_{g}(0,\mu)$), which actually represents the twist-four three-parton correlation [23, 9]. Indeed, extending the treatment of the “virial theorem” in [17] based on the dilatation transformation, we obtain the exact relation,
$\bar{C}_{q}(0,\mu)=\frac{1}{3M}\Bigl\langle\int{{d^{3}}}xx_{\beta}\bar{\psi}(x)\gamma_{\alpha}\,g{F^{\alpha\beta}}(x)\psi(x)\Bigr\rangle\ ,$ (16)
corresponding to the forces between the quarks and gluons, whose details will be discussed elsewhere.
Acknowledgments—The author thanks Hiroyuki Kawamura and Yuichiro Kiyo for insightful discussions. This work was supported by JSPS KAKENHI Grant Numbers JP24K07055 and JP23K03419.

### v2

@@
 The invariant mass of (2) leads to the well-known proton mass formula [21] associated with the trace anomaly [6, 7, 8], i.e., $T^{\alpha}_{\ \alpha}=\frac{\beta(g)}{2g}F^{2}+[1+\gamma_{m}(g)]m\bar{\psi}\psi$, with the QCD $\beta$-function $\beta(g)$ and the anomalous dimension for the quark mass, $\gamma_{m}(g)$.
-Thus, the proton mass is expressed by the non-perturbative matrix elements $\langle p|F^{2}|p\rangle$, $\langle p|m\bar{\psi}\psi|p\rangle$; these matrix elements of the Lorentz scalar operators equal the corresponding matrix elements at the rest frame, as $\langle p|{\cal O}(0)|p\rangle$ $\langle p|{\cal O}(0)|p\rangle|_{\bm{p}=\bm{0}}=2M\langle\int d^{3}x{\cal O}(x)\rangle$, with ${\cal O}(x)=F^{2}(x),m\bar{\psi}\psi(x)$, where we have introduced the shorthand notation for the expectation value in the rest frame as $\langle\,\cdots\,\rangle\equiv(\langle p|\cdots|p\rangle/\langle p|p\rangle)_{\bm{p}=\bm{0}}$.
+Thus, the proton mass is expressed by the non-perturbative matrix elements $\langle p|F^{2}|p\rangle$, $\langle p|m\bar{\psi}\psi|p\rangle$; these matrix elements of the Lorentz scalar operators equal the corresponding matrix elements at the rest frame, as $\langle p|O(0)|p\rangle$ $\langle p|O(0)|p\rangle|_{\bm{p}=\bm{0}}=2M\langle\int d^{3}xO(x)\rangle$, with $O(x)=F^{2}(x),m\bar{\psi}\psi(x)$, where we have introduced the shorthand notation for the expectation value in the rest frame as $\langle\,\cdots\,\rangle\equiv(\langle p|\cdots|p\rangle/\langle p|p\rangle)_{\bm{p}=\bm{0}}$.
 As discussed below, $\langle p|m\bar{\psi}\psi|p\rangle$ is expressed by the sigma terms determined (semi)emprically; then, the condition that the second equation of (2) should reproduce the value of the proton mass allows us to evaluate the proton matrix element $\langle p|F^{2}|p\rangle$.
@@
 $T_{q\ \alpha}^{\alpha}=m\bar{\psi}\psi+\frac{{{\alpha_{s}}}}{{4\pi}}\left(\frac{{{n_{f}}}}{3}F^{2}+\frac{{4{C_{F}}}}{3}m\bar{\psi}\psi\right)+\cdots,$ (3)
-and $T_{g\ \alpha}^{\alpha}=T^{\alpha}_{\ \alpha}-T_{q\ \alpha}^{\alpha}$, with $\alpha_{s}\equiv g^{2}/4\pi$, for $N_{c}$ color with $C_{F}=(N_{c}^{2}-1)/2N_{c}$; (3) holds in the MS-like schemes, and the ellipses stand for the corrections at the two loops (${\cal O}(\alpha_{s}^{2})$) and higher, which are derived in [5, 9, 10].
+and $T_{g\ \alpha}^{\alpha}=T^{\alpha}_{\ \alpha}-T_{q\ \alpha}^{\alpha}$, with $\alpha_{s}\equiv g^{2}/4\pi$, for $n_{f}$ flavor and $N_{c}$ color with $C_{F}=(N_{c}^{2}-1)/2N_{c}$; (3) holds in the MS-like schemes, and the ellipses stand for the corrections at the two loops (${\cal O}(\alpha_{s}^{2})$) and higher, which are derived in [5, 9, 10].
 The quantitative results using the three-loop formulas [9] for (3) will be discussed in Figs.
@@
 Physically, the gravitational form factors represent the coupling of the proton with a graviton.
-Instead of probing them via such impractical graviton-proton processes, the gravitational form factors can be accessed through the spin-2 contributions induced by the two vector currents; e.g., with the two electromagnetic currents, the deeply virtual Compton scatterings [26] in JLab [27] and the future electron-ion collider (EIC) allow us to probe the gravitational form factors.
+Instead of probing them via such impractical graviton-proton processes, the gravitational form factors can be accessed through the spin-2 contributions induced by the two vector currents; e.g., with the two electromagnetic currents, the deeply virtual Compton scatterings [26] in JLab [27] and the future electron-ion collider (EIC) [28] allow us to probe the gravitational form factors.
 Substituting the forward limit of (5) into $M=\langle\,H\,\rangle$, we obtain the “two-term decomposition” [13] using the $\Delta=0$ values of the gravitational form factors as
@@
 $\!\!\!\!\left.A_{q}\left(0,\mu_{0}\right)\right|_{{n_{f}}=3}=0.57(1),\ \left.A_{g}\left(0,\mu_{0}\right)\right|_{{n_{f}}=3}=0.43(1).$ (7)
-The values of $\bar{C}_{r}(0,\mu)$ are also available with the same level accuracy as in $A_{r}(0,\mu)$: the trace of both sides of (5) allows us to write down the formula [5],
+The values of $\bar{C}_{r}(0,\mu)$ are also available with the same level of accuracy as in $A_{r}(0,\mu)$: the trace of both sides of (5) allows us to write down the formula [5],
 $\bar{C}_{r}(0,\mu)=-\frac{1}{4}A_{r}(0,\mu)+\frac{1}{8M^{2}}\langle p|T_{r\ \alpha}^{\alpha}|p\rangle,$ (8)
@@
 For the $n_{f}=3$ active quark flavors, $M_{m}$ of (10) can be expressed as $M_{m}=\langle p|[(m_{u}+m_{d})/2](\bar{u}u+\bar{d}d)+m_{s}\bar{s}s|p\rangle/2M=\sigma_{\pi N}+\sigma_{s}$, up to small isospin-violating corrections of ${\cal O}(m_{d}-m_{u})$, using the pion-nucleon sigma-term $\sigma_{\pi N}$ and the strangeness content $\sigma_{s}$.
-Here, we use $\sigma_{\pi N}=0.0591(35)$ GeV due to a recent phenomenological analysis [29] and $\sigma_{s}=0.0456(62)$ GeV due to a recent lattice QCD determination [30]; these non-perturbative matrix elements arise also in (8) with (3), and the above-mentioned NNLO values of $\bar{C}_{q,g}(0,\mu)$ (e.g., (9)) are calculated using the same input values for $\sigma_{\pi N},\sigma_{s}$.
+Here, we use $\sigma_{\pi N}=0.0591(35)$ GeV due to a recent phenomenological analysis [30] and $\sigma_{s}=0.0456(62)$ GeV due to a recent lattice QCD determination [31]; see also [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
+These non-perturbative matrix elements arise also in (8) with (3), and the above-mentioned NNLO values of $\bar{C}_{q,g}(0,\mu)$ (e.g., (9)) are calculated using the same input values for $\sigma_{\pi N},\sigma_{s}$.
 Our result for (12) using the above-explained non-perturbative inputs provides the NNLO update for the evaluation of Ji’s four-term decomposition of the proton mass, as shown in Fig. 1(a).
@@
 Comparing this with (5), $A_{r}(0,\mu)$ is of twist two, and this implies (iii) above.
-The studies of hard processes like the deep inelastic lepton-proton scatterings (DIS) clarified [19] that the twist-two operators in QCD are associated with the independent partonic motion inside the proton, according to the asymptotic freedom, while the non-perturbative QCD interactions induce the correlations among the partons, emerging as matrix elements of the operators of twist four and higher  [^6] For the quantities dependent on the polarization, the twist-three operators also play roles to describe the spin-dependent partonic correlations, see e.g., [19, 34, 35, 36]..
+The studies of hard processes like the deep inelastic lepton-proton scatterings clarified [19] that the twist-two operators in QCD are associated with the independent partonic motion inside the proton, according to the asymptotic freedom, while the non-perturbative QCD interactions induce the correlations among the partons, emerging as matrix elements of the operators of twist four and higher  [^6] For the quantities dependent on the polarization, the twist-three operators also play roles to describe the spin-dependent partonic correlations, see e.g., [19, 46, 47, 48]..
 As noted in (iii), $M_{q}$ does not match with such physical meaning of the twist, although $M_{q}$ is intended [11, 4] to represent the contribution of the quark’s gauge-invariant “kinetic and potential energies” operator $\bar{\psi}i\bm{\gamma}\cdot\overleftrightarrow{\bm{D}}\psi$ of (4).
@@
 Figure 3: Same as Fig. 1, but for the pion case.
-As explained in [20], the global QCD analyses for the pion’s PDFs are available at the NLO level, and the resultant value, $A_{q}^{\pi}(0,\mu_{0})=0.58(9)$ [40], has rather large ($\sim$ ten percent level) uncertainties (see also [41, 42, 43]), while chiral perturbation theory gives the accurate value for $M_{m}^{\pi}=\langle\pi|m\bar{\psi}\psi|\pi\rangle/2M^{\pi}$ corresponding to (10): $M_{m}^{\pi}=M^{\pi}/2$, to the $O(6\%)$ correction [44, 45].
-Using these input values, $A_{q,g}^{\pi}(0,\mu)$, $\bar{C}^{\pi}_{q,g}(0,\mu)$ have been calculated [20] by our NNLO formulas in the $\overline{\rm MS}$ scheme as, e.g., $\bar{C}^{\pi}_{q}(0,\mu_{0})|_{{n_{f}}=3}=-\bar{C}^{\pi}_{g}(0,\mu_{0})|_{{n_{f}}=3}=-0.03(2)$ with $\mu_{0}=2$ GeV.
+As explained in [20], the global QCD analyses for the pion’s PDFs are available at the NLO level, and the resultant value, $A_{q}^{\pi}(0,\mu_{0})=0.58(9)$ [52], has rather large ($\sim$ ten percent level) uncertainties (see also [53, 54, 55]), while chiral perturbation theory gives the accurate value for $M_{m}^{\pi}=\langle\pi|m\bar{\psi}\psi|\pi\rangle/2M^{\pi}$ corresponding to (10): $M_{m}^{\pi}=M^{\pi}/2$, to the ${\cal O}(6\%)$ correction [56, 57].
+Using these input values, $A_{r}^{\pi}(0,\mu)$, $\bar{C}^{\pi}_{r}(0,\mu)$ have been calculated [20] by our NNLO formulas in the $\overline{\rm MS}$ scheme as, e.g., $\bar{C}^{\pi}_{q}(0,\mu_{0})|_{{n_{f}}=3}=-\bar{C}^{\pi}_{g}(0,\mu_{0})|_{{n_{f}}=3}=-0.03(2)$ with $\mu_{0}=2$ GeV.
 These values allow us to evaluate the pion-mass decompositions corresponding to the four-term formulas (12) and (13), as displayed in Figs.